A solution of sodium citrate (monobasic) is prepared to a formal concentration of 0.2 M. If the pKa's for the triprotic citric acid molecule are 3.13, 4.76, 6.40 what is the approximate pH of this solution
Answer #1

Consider equilibira of citric acid as shown below.

H3Cit H + + H2Cit -   H + + HCit 2- H + + Cit 3-

Sodium citrate mono basic is a intermediate form of Citric acid and disodium citrate.

We know that, for intermediate form, [H+ ] = K 1 K 2 F + K 1 K w / K 1 + F

We have , pK 1 = 3.13

We know that, pK 1 = -log K 1 . Therefore, K 1 = 10 - pK 1 = 10 - 3.13 = 7.413 10 -04

Similarly , K 2 = 10 - pK 2 = 10 - 4.76 = 1.738 10 -05

Here F = 0.2 M and K w = 1 10 -14

[H+ ] = (7.413 10 -04) (1.738 10 -05 ) 0.2 + (7.413 10 -04 1  10 -14 ) / 7.413 10 -04 + 0.2

= 2.577 10 -09 / 0.20074

= 1.284 10 -08

= 1.133 10 -04 M

We have, pH = - log [H+ ] = - log 1.133 10 -04 = 3.946

ANSWER : Approximate pH of 0.2 M solution of mono sodium citrate solution = 3.946